OpenStudy (anonymous):

Solve the equation by factoring. x^2+5x=6

OpenStudy (anonymous):

\[x^{2}+5x-6=x^{2}+6x-x-6=x(x+6)-1(x+6) = (x+6) (x-1)\]

OpenStudy (anonymous):

x=-6 and 1

OpenStudy (anonymous):

Still kind of confusing

OpenStudy (anonymous):

which part?

OpenStudy (zale101):

@crazysingh got x^2+5x-6=0 by subtracting +6 to both sides x^2+5x=6 x^2+5x-6=6-6 x^2+5x-6=0 just in case if you got confused by this part

OpenStudy (anonymous):

Can you maybe show me step by step?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you are given \[x^{2}+5x=6\] subtract 6 from both sides.

OpenStudy (anonymous):

you will get \[x ^{2}+5x-6=0\]

OpenStudy (anonymous):

now I can write \[5x=6x-x\]

OpenStudy (anonymous):

so the equation becomes \[x ^{2}+6x-x-6=0\]

OpenStudy (anonymous):

are you getting me?

OpenStudy (anonymous):

How'd you get 5x if you subtracted 6 from both sides?

OpenStudy (anonymous):

because its already there :/

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

So wait...which two numbers are you subtracting 6 from?